OpenStudy (anonymous):

what is y in the following system x+3y =6 2x-7= 10

OpenStudy (anonymous):

@dan815 @whpalmer4

OpenStudy (anonymous):

pleasee

OpenStudy (unklerhaukus):

solve the second equation for x, then substitute that into the first

OpenStudy (anonymous):

This guy knows what he is talking about ^

OpenStudy (anonymous):

oh whoops that is a typo. the 7 is supposed to be Y

OpenStudy (anonymous):

2x-y= 10

OpenStudy (anonymous):

so what do i do then ?

OpenStudy (unklerhaukus):

x+3y = 6 2x -y = 10 multiply the second equation by three then add it to the first equation the y terms will drop out, so you can solve for x then substitute this x into the first equation, and you'll be able to solve for y

OpenStudy (anonymous):

x = 5.142 ?????

OpenStudy (whpalmer4):

That's the elimination approach. Another approach, for which this problem is better suited than some, is substitution. The first equation is easy to solve for x. x = 6-3y. Now substitute (6-3y) in the second equation wherever you see x. That gives you an equation only in terms of y. Solve that. Plug the value into x = 6-3y to get the value of x. You should get the same result either way; sometimes one approach is easier to do the work than the other.

OpenStudy (whpalmer4):

Yes, that's correct, but I would leave the answer in fraction form: x = 36/7 unless someone specifically asks for a decimal.

OpenStudy (anonymous):

how would i plug that in?

OpenStudy (anonymous):

36/7 + 3y = 6

OpenStudy (whpalmer4):

how would you plug what in? x=36/7? 36/7 + 3y = 6 multiply both sides by 7 to get rid of the fraction 36/7*7 + 7*3y = 7*6 36 + 21y = 42 you can do the rest, right?

OpenStudy (anonymous):

oooohhhh

OpenStudy (anonymous):

i got y = .28

OpenStudy (anonymous):

im so lost -_-

OpenStudy (whpalmer4):

y = 2/7, yeah

OpenStudy (anonymous):

the answer choices are in slope intercept form

OpenStudy (anonymous):

i think 2/7 can be 1/3 .

OpenStudy (anonymous):

choices are : 1/3x + 2 -1/3 x+6 -x+2 -1/3x + 2

OpenStudy (whpalmer4):

are you sure you haven't mixed two problems together? x+3y=6 solve for y 3y = 6-x y = 2-x/3 y = (-1/3)x + 2

OpenStudy (anonymous):

hm. maybe i did im really tired. can you help with some more

OpenStudy (whpalmer4):

we can find out :-)